L-shaped room consisting of a rectangle that is 8 1/2 ft by 11ft and one that is 6 1/2ft 7 1/2 ft
-what is the area
-what is the perimeter?

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

figure area for each section then total
8.5*11=93.5 sqft
7.5*6.5=48.75 sqft
93.5+48.75=142.25 sq ft total
------------------------------
P=8.5+2+6.5+2(7.5)+2(11)
P=54 ft

Answer by richwmiller(17219)

Are the two rectangles overlapping?